OpenStudy (randomecorner):

What is when z = −3?

TheSmartOne (thesmartone):

What is this mysterious space supposed to equal?

TheSmartOne (thesmartone):

Copy-paste error.

TheSmartOne (thesmartone):

What is what???

OpenStudy (randomecorner):

srry

OpenStudy (randomecorner):

@TheSmartOne

TheSmartOne (thesmartone):

Plug in z=-3 and solve so \(\sf\Large\frac{4(-3)-3}{-3+6}\)

OpenStudy (randomecorner):

so 4 over nine?

OpenStudy (randomecorner):

no?

OpenStudy (randomecorner):

ummm help?

TheSmartOne (thesmartone):

What is \(\sf 4 \times (-3)=?\) What is \(\sf (-3)+6=?\)